OpenStudy (deni):

Solve 0.25(7+3m)=-0.25m

OpenStudy (helpwithk12):

hi

OpenStudy (deni):

Hi

OpenStudy (3mar):

Hi

OpenStudy (helpwithk12):

-1.75 is the answer

OpenStudy (3mar):

It is not useful to just give you the answer without you choose it and know why!

OpenStudy (helpwithk12):

sorry i mean m

OpenStudy (3mar):

1st: solve for m

OpenStudy (helpwithk12):

peace give medal if helped

OpenStudy (3mar):

@Deni

OpenStudy (mathmale):

Start over! "Solve 0.25(7+3m)=-0.25m" First, you could simplify this problem by dividing both sides by 0.25. Doing that will remove the decimal fractions and leave you with the much simpler 7+3m=-m. All you have to do now is to solve 7+3m=-m. Group all "m" terms together on one side of your equation. Show your work if you want feedback on it.